OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -9, an = 4 • an-1

OpenStudy (fibonaccichick666):

so any ideas on how to get \(a_2\)?

OpenStudy (cwrw238):

first term is -9 and subsequent terms are obtained by multiplying by ?

OpenStudy (anonymous):

oh, i think I understand. then i would do 4 * -9 and get -36. then 4*-36 and so on.

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

can you help me on another problem?

OpenStudy (anonymous):

if you dont mind..

OpenStudy (anonymous):

Find the derivative of f(x) = negative 8 divided by x at x = 11

OpenStudy (cwrw238):

f(x) = -8/x = -8 x^-1 f'(x) = 8^x^-2 or 8 / x*2 just plug in x=11 and you have it

OpenStudy (anonymous):

wait where do i plug it in at?

OpenStudy (cwrw238):

8 / x^2 = 8 / 11^2

OpenStudy (cwrw238):

= 8/121

OpenStudy (anonymous):

omg thank you

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

are you busy with something else? i need help with one more question if you dont mind